# A New Era in Computational Pathology: A Survey on Foundation and Vision-Language Models

### v1

@@ IV Vision-Language Models
First, different categorizations of VLMs are provided to give insight into how VLMs are utilized to solve pathology-specific tasks. Second, the common architectural components in VLMs and adopted strategies are summarized. Lastly, a brief overview of models that do not solve a direct pathology task but perform other types of vision-language tasks is summarized.
@@ IV-B Architectural components for VLMs in CPath
IV-C Related works for VLMs in CPath
There are several research works that do not directly utilize VLMs to solve a task in the pathology domain, but investigate or use VLMs for other reasons. One example is Thota et al  [^5] https://github.com/jaiprakash1824/VLM_Adv_Attack [201] which investigates the effect of Projected Gradient Descent (PGD) adversarial perturbation attack on PLIP [40] architecture. Another work by Lucassen et al. [^6] https://github.com/RTLucassen/report_preprocessing [202] provides a pathology report processing workflow that can be used for VLMs in CPath.
Another category of VLMs in CPath that does not directly solve pathology-specific tasks is text-guided diffusion models for image generation. However, these models can perform tasks like virtual stain transfer which can later be used to solve a pathology-specific task. Another use case is extending dataset size with generated synthetic images. One such work is PathLDM [^7] https://github.com/cvlab-stonybrook/PathLDM [203], that performs text-to-image generation on the TCGA-BRCA dataset.

### v2

@@ IV Vision-Language Models
First, different categorizations of VLMs are provided (section IV-A) to give insight into how VLMs are utilized to solve pathology-specific tasks. Second, the common architectural components in VLMs and adopted strategies are summarized (section IV-B). Then, brief summary of models focusing on visual instruction tuning is provided (section IV-C). Lastly, a brief overview is given for models that do not solve a direct pathology-specific task but perform other types of vision-language tasks (section IV-D).
@@ IV-B Architectural components for VLMs in CPath
IV-C Visual Instruction Tuning in CPath
There are four research works/models that fall in this category (PA-LLaVA [163], PathInsight [164], CLOVER [165], PathGen-LLaVA [51] and Quilt-LLaVA [56]).
In PA-LLaVA, a domain-specific large language-vision assistant is proposed which is a LLaVA-based model. In terms of architecture, PLIP is used as the vision encoder and LLama3 is used as the LLM (in conjunction with LoRA [124]). A connector module that consists of self-attention and cross-attention blocks is leveraged to project image embeddings to match the dimension of language tokens. PathInsight leverages three different multi-modal models LLaVA, Qwen-VL [166] and InternLM [167] (in conjunction with LoRA or full-parameter tuning). As the vision encoder either ViT or CLIP is used. CLOVER [165] proposes a framework for cost-effective instruction learning in CPath. It utilizes Quilt-1M as the vision-language dataset, BLIP-2 [138] for vision-language pre-training and a frozen visual encoder (EVA-ViT-G/14 [168]) and a frozen LLM (FlanT5 [116] or Vicuna [54]). The next work PathGen-LLaVA replaces the vision encoder for the CLIP model in LLaVA-v1.5 with PathGen-CLIP and as the frozen LLM Vicuna is used. The last work in this category Quilt-LLaVA is initialized with the general-domain LLaVA. Then it is tuned with QUILT for histopathological domain-alignment and finally, Quilt-instruct is used for visual instruction tuning.
IV-D Related works for VLMs in CPath
There are several research works that do not directly utilize VLMs to solve a task in the pathology domain, but investigate or use VLMs for other reasons. One example is Thota et al [169] which investigates the effect of projected gradient descent (PGD) adversarial perturbation attack on PLIP [42] architecture. Another work by Lucassen et al. [170] provides a pathology report processing workflow that can be used for VLMs in CPath.
Another category of VLMs in CPath that does not directly solve pathology-specific tasks is text-guided diffusion models for image generation. However, these models can perform tasks like virtual stain transfer which can later be used to solve a pathology-specific task. Another use case is extending dataset size with generated synthetic images. One such work is PathLDM [171], that performs text-to-image generation on the TCGA-BRCA dataset.